# CBMC-V3: A CNS-inspired Control Framework Towards Manipulation Agility with SNN

### v2

@@ Abstract
As robotic arm applications extend beyond industrial settings into service-oriented sectors such as catering, household and retail, existing control algorithms struggle to achieve the agile manipulation required for complex environments with dynamic trajectories, unpredictable interactions, and diverse objects. This paper presents a biomimetic control framework based on Spiking Neural Networks (SNN), inspired by the human Central Nervous System (CNS), to achieve agile control in such environments. The proposed framework features five control modules (cerebral cortex, cerebellum, thalamus, brainstem, and spinal cord), three hierarchical control levels (first-order, second-order, and third-order), and two information pathways (ascending and descending). Each module is fully implemented using SNN. The spinal cord module uses spike encoding and Leaky Integrate-and-Fire (LIF) neurons for feedback control. The brainstem module employs a network of LIF and non-spiking LIF neurons to dynamically adjust spinal cord parameters via reinforcement learning. The thalamus module similarly employs a network of LIF and non-spiking LIF neurons to adjusts the cerebellum’s torque outputs via reinforcement learning. The cerebellum module, which provides feedforward gravity compensation torques, uses a recurrent SNN to learn the robotic arm’s dynamics through regression. The framework is validated both in simulation and on real-world robotic arm platform under various loads and trajectories. Results demonstrate that our method outperforms the industrial-grade position control in manipulation agility.
@@ I Introduction
In recent years, robotic arms are increasingly applied in service-oriented sectors such as catering, household and retail. Compared with traditional industrial settings, these scenarios feature more complex environments and diverse tasks while requiring better agility, thereby posing new challenges to control strategies. For instance, a 7-DOF robotic arm can readily cut a circular contour on a workpiece or transport heavy replacement parts, yet it still struggles to perform a task as subtle as peeling an apple, which can be easily accomplished by the human arm. Compared with biological systems, even the most advanced robotic arms today fall short of the agility of human upper limb [mason2018], underscoring the existence of advanced motor control mechanisms in humans that remain underutilized in robotic systems.
In well-modeled environments, control algorithms combining Whole-Body Control (WBC) [Abe2007, liu2011interactive, weighttans2011, Lee2012Intermediate, Escande2014, Kim2019, han2021recursive] and Model Predictive Control (MPC) [carron2019data, pankert2020perceptive, wang2023trajectory] can effectively counteract disturbances and achieve strong control performance. Such model-based approaches reflect human capabilities in applying prior knowledge to solve specific tasks. However, they lack the adaptability inherent to human motor learning. In complex scenarios, variations in the size, mass, or materials of manipulated objects may render the kinematic and dynamic models inaccurate, limiting the effectiveness of these model-based strategies.
@@
With the rapid progress of neural networks, their powerful approximation and learning capabilities have motivated researchers to employ them for adaptive model correction in response to environmental changes. For example, [chaoui2009ann] designed feedforward and feedback neural networks to compensate for joint flexibility and friction; [he2017adaptive] employed fuzzy neural networks with adaptive impedance control to handle unknown dynamics and state constraints; [wang2018adaptive] proposed a neural-network-based adaptive controller capable of addressing nonlinear dynamics and external disturbances without prior models; [salloom2020adaptive] integrated genetic algorithm observers with neural networks to compensate for underwater disturbances; [liu2019adaptive] investigated the role of hidden-layer size in adaptive control and proposed dynamic adjustment strategies; and [pham2020adaptive] applied radial basis function networks with dynamic surface control for robust dual-arm coordination. Despite these progression, they rely on artificial neural networks (ANN), which inherently disregard the temporal-spatial characteristics of motion, focusing only on instantaneous control. Furthermore, they remain constrained by robotic dynamics models, with network structures and training methods often tailored to specific problems and tightly coupled to complex mathematical formulations, thereby limiting generalizability.
Recently, spiking neural networks (SNN) combined with brain-inspired architectures have emerged as promising alternatives. By emulating biological neurons, SNNs process and store information through spike trains and membrane potential dynamics, offering advantages in computational speed, energy efficiency, information dimension, and biological plausibility, making them well-suited for robotic control [bing2018survey]. For instance, [carrillo2008real] developed a cerebellum-inspired SNN trained via spike-timing-dependent plasticity (STDP) to learn manipulator dynamics for torque control; [abadia2019robot] extended this model to human–robot collaboration scenarios, demonstrating robustness under unknown disturbances; [abadia2021cerebellar] further established resilience to control delays, highlighting potential in teleoperation and cloud robotics; [doi:10.1126/scirobotics.adp2356] refines this framework to enable robotic arms to dynamically adjust compliance. However, these methods suffer from structural complexity, computational cost, hyperparameter sensitivity, and limited interpretability, as they focus narrowly on cerebellar functionality while neglecting the integrative role of the broader central nervous system (CNS), which coordinates control through interactions among several neural areas[Halassa_2022].
Inspired by this systemic perspective of the CNS, [li2023cbmc] proposed the CBMC framework, incorporating modules such as the cortex, cerebellum, and spinal cord. Yet its computational inefficiency restricted real-time applicability. Building on this, [10769842] introduced a hierarchical control architecture deployable on neuromorphic hardware, though its lack of neural networks in some modules compromised accuracy.
Despite these advances, current brain-inspired algorithms have yet to demonstrate agile performance comparable to that of humans. As defined in [8610238], agility is a capability characterized by rapid and efficient transitions between diverse motion patterns, with precision and speed as its key attributes. In this work, we propose CBMC-V3, a fully SNN-based brain-inspired control framework designed to achieve agile robotic arm control. The key contributions are fourfold: (1) Proposing a CNS-inspired control framework leveraging hierarchical modularity; (2) Designing a fully SNN-based implementation; (3) Validating the framework and module design in simulation experiments; (4) Deploying the algorithm on a robotic platform, demonstrating agility.
@@
The remainder of this paper is organized as follows: Section II introduces the framework structure; Section III details the SNN-based implementations of individual modules; Section IV presents simulation analyses; Section V conducts experiments on a robotic arm platform, and Section VI concludes the study.
Figure 1: Overview of the human CNS and the proposed framework structure. (a) Human CNS relating to motion control. Sensory information and motor commands are transmitted through three orders of neurons along ascending and descending pathways. Sensory signals travel from peripheral receptors to the spinal cord, brainstem, and ultimately the cerebellum and cerebral cortex, while motor commands descend from the cerebral cortex and cerebellum through the thalamus and spinal cord to the muscles. The spinal cord can independently generate reflexive responses via reflex arcs, enabling rapid reactions to stimuli [LLOBERA2023190]. The brainstem contributes to movement termination, reorientation, and coordination with the spinal cord [merel2019hierarchical, annurev:/content/journals/10.1146/annurev-neuro-082321-025137]. The thalamus functions as a hub for motor coordination, multimodal perception, and sensorimotor integration [LLOBERA2023190, Halassa_2022]. The cerebellum regulates body movements, balance, and posture by integrating sensory inputs and refining motor commands [LLOBERA2023190]. The cerebral cortex generates decisions and motion trajectories. (b) Schematic diagram of our proposed framework, featuring five modules, three control levels and two pathways as mentioned in (a). (c) Neural network design of each module, utilizing three types of spiking neurons: input fiber (grey), LIF neuron (yellow) and non-spiking LIF neuron (green).
@@ II Framework structure
Inspired by the motion control mechanism of the human CNS, as illustrated in Fig.1a, the structure of our proposed framework CBMC-V3 is shown in Fig.1b. This framework consists of three fundamental components: (1) five functional modules, namely the cerebral cortex module, cerebellum module, thalamus module, brainstem module, and spinal cord module; (2) three hierarchical control levels, namely the first-order, second-order, and third-order level; (3) two information pathways, the ascending and descending pathway. In the following sub-sections, we will introduce each of these components in detail and explain how the framework integrates functional modules, control loops, and information pathways to accomplish a complete operational cycle.
II-A Module function
@@
In the human CNS, the spinal cord can generate reflexive actions independently via reflex arcs, enabling rapid responses to sensory stimuli without cortical input [LLOBERA2023190]. Inspired by this mechanism, the spinal cord module in our framework implements a PD-like feedback control function, producing torque commands at real-time control frequencies, i.e.,
@@
The brainstem has been shown to contain pathways involved in movement termination and reorientation, and to interact with the spinal cord to coordinate motor behaviors [merel2019hierarchical, annurev:/content/journals/10.1146/annurev-neuro-082321-025137]. Motivated by this, the brainstem module in our framework dynamically adjusts the PD gains for each joint based on trajectory information, providing the spinal cord module with time-varying proportional and derivative coefficients, i.e.,
$\bm{K}_{V}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})=diag(k_{v,1},k_{v,2},\cdots,k_{v,n})$ (2)
$\bm{K}_{P}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})=diag(k_{p,1},k_{p,2},\cdots,k_{p,n})$ (3)
By integrating spinal feedback control with brainstem-driven adaptive gain modulation, the framework circumvents the complexity of manipulator dynamics while compensating for structured uncertainties through learning.
The gravity compensation torque of a robotic manipulator, denoted as $\bm{\tau}_{g}(\bm{q})$, is a feedforward torque designed to counteract the gravitational forces acting on the manipulator itself. It plays a crucial role in ensuring the smoothness and stability of motion. In human motor control, the cerebellum is primarily responsible for coordinating and regulating body movements, ensuring muscle and postural balance. It integrates sensory inputs, such as muscle tension and joint position, and adjusts the magnitude and direction of motor commands to achieve precise and fluid motion [LLOBERA2023190]. Analogously, gravity compensation in robotic manipulators serves a function comparable to that of the cerebellum, and thus the cerebellar module in our framework is tasked with providing this compensation. Furthermore, we note that during human arm manipulation, accurate models of dynamics are not strictly required. Instead, the brain relies on stored motor patterns to cope with different loading conditions. For instance, when grasping an object, we estimate its weight based on prior experience and adjust the exerted force accordingly. Such experiential knowledge is typically not acquired on site during execution, but rather accumulated over time through long-term daily interactions. Inspired by this biological mechanism, the proposed framework equips the cerebellum module with the ability to pre-learn multiple gravity compensation ‘patterns’ corresponding to different load masses. These patterns can then be retrieved and combined to generate an $n\times m$ set of compensation torques,
as a function of joint positions, where $\tau_{\text{cb},ji}(\bm{q}),j=1,\cdots,n,i=1,\cdots,m$, denotes the torque provided for the $j$-th joint under the $i$-th compensation pattern.
Although suitable motor patterns can often be selected based on prior experience, in certain cases, such selections may be inaccurate due to factors such as visual deception. For instance, the mass of an opaque box cannot be reliably estimated visually, which may lead to insufficient force to move the load or, conversely, an excessive initial force causing large errors. Nevertheless, humans can quickly adjust their motor output upon physically contacting the object, using sensory feedback to correct their movement patterns. This phenomenon is closely associated with the function of the thalamus, which serves as a central hub for motor coordination and plays a critical role in multimodal perception and sensorimotor integration [LLOBERA2023190, Halassa_2022]. Analogously, a robotic arm can ‘perceive’ the load weight via force sensors installed at the end-effector. The thalamus module, similar to humans, uses the sensed load information to dynamically weight and integrate the torque contributions provided by the cerebellar module, generating an actual gravity compensation torque, i.e.,
where $\bm{F}_{\text{ef}}$ denotes the end-effector torque sensor feedback, $\bm{w}(\bm{F}_{\text{ef}})=(w_{1},w_{2},\dots,w_{m})^{T}$ represents the dynamically learned weight coefficients of the thalamus module, and $\tau_{g,1},\dots,\tau_{g,n}$ are the resulting gravity compensation torques. By combining the prior knowledge stored in the cerebellum module with the adaptive weighting capability of the thalamus module, the framework effectively addresses the unknown dynamics of the manipulator’s end-effector load.
The resulting torque is then summed with the feedback torque generated by the spinal cord module and applied as joint-level control commands to the actuators, i.e.,
$\bm{\tau}=\bm{\tau}_{\text{fb}}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})+\bm{\tau}_{g}(\bm{q},\bm{F}_{\text{ef}})=(\tau_{1},\cdots,\tau_{n})^{T}$ (6)
where $\bm{\tau}_{\text{fb}}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})$ is produced by the spinal cord and brainstem module, $\bm{\tau}_{g}(\bm{q},\bm{F}_{\text{ef}})$ is produced by the cerebellum and thalamus module, and $(\tau_{1},\dots,\tau_{n})^{T}$ represents the joint torques delivered to the actuators.
In conclusion, the function of each module can be summarized as follows:
  • The cerebral cortex module makes decisions and generates motion trajectories. In this research, it stores and replays pre-planned trajectories.
  • The cerebellum module is a cerebellar-like network pre-trained to store several fundamental gravity compensation motion patterns.
  • The thalamus module learns to generate compensation torque adaptable to varying loads by online RL using proprioception information and the cerebellum module output.
  • The brainstem module learns to provide appropriate feedback control coefficients by online RL using proprioception information.
  • The spinal cord module combines feedback control coefficients from the brainstem and torque from the thalamus module to generate real-time control torque commands.
II-B Hierarchical control level

### v3

# CBMC-V3: A CNS-inspired Control Framework Towards Agile Manipulation with SNN

@@ Abstract
As robotic arm applications expand beyond traditional industrial settings into service-oriented domains such as catering, household, and retail, existing control algorithms struggle to achieve the level of agile manipulation required in unstructured environments characterized by dynamic trajectories, unpredictable interactions, and diverse objects. This paper presents a biomimetic control framework based on spiking neural network (SNN), inspired by the human central nervous system (CNS), to address these challenges. The proposed framework comprises five control modules—cerebral cortex, cerebellum, thalamus, brainstem, and spinal cord—organized into three hierarchical control levels (first-order, second-order, and third-order) and two information pathways (ascending and descending). All modules are fully implemented using SNN. The framework is validated through both simulation and experiments on a commercial robotic arm platform across a range of control tasks. The results demonstrate that the proposed method outperforms the baseline in terms of agile motion control capability, offering a practical and effective solution for achieving agile manipulation.
@@ I Introduction
In recent years, robotic arms have been increasingly deployed in manipulation-intensive domains such as catering, household, and retail services, posing new challenges for control strategies. While a 7-DOF robotic arm can readily cut precise circular contours or transport heavy replacement parts, it still struggles with delicate tasks such as peeling an apple—an operation that the human arm performs effortlessly. This gap largely arises from humans’ strong motion agility, which enables robust and flexible handling of delicate, contact-rich interactions. However, defining and quantifying agility in motion control remains challenging due to its context-dependent and multifaceted nature.
Drawing insights from [39, 12], we characterize agile motion control using four key features: (1) precision, (2) adaptation, (3) resilience, and (4) energy efficiency. Precision measures an algorithm’s ability to accurately track a desired trajectory, for example in terms of end-effector position and orientation tracking errors. Adaptation evaluates the ability to maintain tracking precision across different tasks, such as under varying load masses and trajectory configurations. Resilience reflects how quickly the algorithm recovers from unstructured interactions or disturbances. Energy efficiency quantifies the computational resources required by the algorithm.
Control algorithms that combine Whole-Body Control (WBC) [4, 27, 36, 24, 13, 23, 19] and Model Predictive Control (MPC) [9, 33, 44] can achieve high precision in well-defined environments and typically exhibit high energy efficiency. However, their resilience is often compromised as model complexity increases. Moreover, these methods lack effective mechanisms to cope with environmental variations and model inaccuracies, such as objects with differing shapes or masses and elastic or biomimetic joints, resulting in poor adaptation.
@@
Artificial neural network (ANN) can learn and adapt to environmental changes, resulting in relatively strong adaptation capabilities. When combined with model-based dynamics, it can also achieve relatively high precision. For example, [10] designed feedforward and feedback neural networks to compensate for joint flexibility and friction; [20] employed fuzzy neural networks with adaptive impedance control to handle unknown dynamics and state constraints; [43] proposed a neural-network-based adaptive controller capable of addressing nonlinear dynamics and external disturbances without prior models; [37] integrated genetic algorithm observers with neural networks to compensate for underwater disturbances; [26] investigated the role of hidden-layer size in adaptive control and proposed dynamic adjustment strategies; and [35] applied radial basis function networks with dynamic surface control for robust dual-arm coordination. However, these methods typically exhibit limited resilience and energy efficiency due to the computational characteristics of ANN [6]. Moreover, the limited ability of ANN to model temporal dynamics further restricts their adaptation in highly dynamic manipulation tasks.
Recently, spiking neural network (SNN) combined with cerebellum-inspired architecture has emerged as promising alternatives. By emulating biological neurons, SNN processes and stores information through spike trains and membrane potential dynamics, offering significant advantages in energy efficiency [6]. [8] developed a cerebellum-inspired SNN trained via spike-timing-dependent plasticity (STDP) to learn robotic arm dynamics for torque control; [2] extended this model to human–robot collaboration scenarios, demonstrating robustness under unknown disturbances; [3] addressed control delays using this network, highlighting its potential for teleoperation and cloud robotics; and [1] further refined the framework to enable robotic arms to dynamically adjust compliance. By modeling human neural processing at the level of individual neurons and closely mimicking the structure of the human cerebellum, these approaches can exhibit human-like behavior and strong adaptation capabilities. However, despite the low per-neuron energy consumption, the extremely large network scale (e.g., over 60k neurons and 32M synaptic connections) limits the overall energy efficiency of the framework. Moreover, the use of unsupervised, biologically inspired learning rules such as STDP—whose convergence properties lack theoretical guarantees—constrains achievable precision and resilience.
In our previous work CBMC-V2 [32], we presented a hierarchical control framework inspired by the central nervous system (CNS). Compared with cerebellum-inspired SNN approaches, this framework achieves higher energy efficiency by modular design and neuromorphic hardware deployment, and attains improved precision and resilience by incorporating gradient-based learning into both the cerebral cortex and cerebellum modules. However, the framework exhibits weak adaptation due to limited learning capability, arising from insufficient sensory inputs (e.g., load information), the absence of SNN in some modules, and the tight coupling between control and planning in the cerebral cortex module. Furthermore, because the cerebellum module largely preserves the network structure adopted in [2, 3, 1], the framework inherits their inherent limitations, which constrain the achievable precision, resilience, and energy efficiency.
Overall, this research aims to provide a motion control framework towards agile manipulation by addressing the limitations identified in CBMC-V2. To this end, we propose CBMC-V3, a CNS-inspired, fully SNN-based control framework. The main contributions of this work are threefold: (1) the proposal of a CNS-inspired control framework; (2) the design of a fully SNN-based implementation; and (3) the demonstration of agile motion control in both simulation and on a commercial robotic arm.
@@
The remainder of this paper is organized as follows: Section II introduces the framework structure; Section III details the SNN-based implementations of individual modules; Section IV presents simulation results; Section V reports experiment results on a commercial robotic arm, and Section VI concludes this paper.
Figure 1: Overview of the human CNS and the proposed framework. (a) Human CNS relating to motor control. (b) Diagram of our proposed framework, featuring five modules, three control levels and two pathways. (c) Neural network design of each module with three types of spiking neurons: connecting fiber (grey), LIF neuron (yellow) and non-spiking LIF neuron (green).
@@ II Framework structure
Inspired by the motor control mechanism of the human CNS, as illustrated in Fig.1a, the structure of our proposed framework CBMC-V3 is shown in Fig.1b. This framework consists of three key aspects: (1) five modules including the cerebral cortex module, the cerebellum module, the thalamus module, the brainstem module, and the spinal cord module; (2) three hierarchical control levels, namely the first-order level, second-order level, and third-order level; (3) two information pathways, i.e. the ascending pathway and the descending pathway. In the following sub-sections, we will introduce each of these aspects and present the pseudocode implementation of this framework.
II-A Modules
II-A1 The cerebral cortex module
The cerebral cortex module makes decisions and generates motion trajectories. In this research, it stores and replays pre-planned velocity $\bm{\dot{q}}_{d}\in\mathbb{R}^{n}$ and position $\bm{q}_{d}\in\mathbb{R}^{n}$ of the $n$ actuated joints.
II-A2 The cerebellum module
The cerebellum module is a cerebellar-like network pre-trained to store gravity compensation motor patterns for corresponding load masses. In human motor control, the cerebellum is responsible for coordinating and regulating body movements, ensuring muscle and postural balance. It integrates sensory inputs, such as muscle tension and joint position, and adjusts the magnitude and direction of motor commands to achieve precise and fluid motion [28]. The gravity compensation torque for a robotic arm is designed to counteract the gravitational forces acting on the arm itself. It plays a crucial role in ensuring the smoothness and stability of motion. The cerebellum can be regarded as performing a function similar to gravity compensation, and the cerebellum module in our framework is therefore designed to provide this compensation.
Furthermore, we notice that during human manipulation, accurate dynamic model is not required. Instead, the brain relies on stored motor patterns to cope with different load conditions. For instance, when grasping an object, we estimate its mass based on prior experience and adjust the exerted force accordingly. Such knowledge is typically not acquired on site during execution, but rather accumulated over time through long-term daily interactions. Inspired by this biological mechanism, we equip the cerebellum module with the ability to pre-learn $m$ gravity compensation patterns corresponding to different load masses. These patterns can then be retrieved and combined in prediction to generate a compensation torque matrix
as a function of actual joint positions $\bm{q}\in\mathbb{R}^{n}$, where $\tau_{\text{cb},ji}(\bm{q}),j=1,\cdots,n,i=1,\cdots,m$ denotes the torque provided for the $j$-th joint under the $i$-th compensation pattern.
II-A3 The thalamus module
The thalamus module learns to generate compensation torque adaptable to arbitrary loads by online RL using proprioceptive information and the cerebellum module output. In unstructured environments, the selection of motor patterns may be inaccurate due to factors such as visual deception. For instance, the mass of an opaque box cannot be reliably estimated visually, which may lead to insufficient force to move the load or, conversely, an excessive initial force causing large errors. Nevertheless, humans can quickly adjust motor output upon contacting the object, using sensory feedback to correct their movement patterns. This phenomenon is closely associated with the function of the thalamus, which serves as a central hub for motor coordination and plays a critical role in multimodal perception and sensorimotor integration [28, 18].
Analogously, a robotic arm can perceive the load mass via force sensors installed at the end-effector. The thalamus module, similar to humans, uses the sensed load information to appropriately integrate the torque matrix provided by the cerebellum module, i.e.,
where $\bm{F}_{\text{ef}}$ denotes the end-effector torque sensor feedback, $\bm{w}(\bm{F}_{\text{ef}})=(w_{1},w_{2},\dots,w_{m})^{\top}$ represents the $m$ dynamically learned weight coefficients of the thalamus module, and $\bm{\tau}_{g}(\bm{q},\bm{F}_{\text{ef}})=(\tau_{g,1},\dots,\tau_{g,n})^{\top}$ are the resulting gravity compensation torque of each joint for arbitrary load mass.
By combining the prior knowledge stored in the cerebellum module with the adaptive weighting capability of the thalamus module, the framework effectively addresses the unknown dynamics of the robotic arm’s end-effector load.
II-A4 The brainstem module
The brainstem module learns to provide appropriate feedback control coefficients by online RL using proprioceptive information. The brainstem has been shown to contain pathways involved in movement termination and reorientation, and to interact with the spinal cord to coordinate motor behaviors [29, 25]. Motivated by this, the brainstem module in our framework dynamically adjusts the proportional and derivative coefficients of the $n$ actuated joints based on trajectory information, i.e.,
$diag(k_{p,1},k_{p,2},\cdots,k_{p,n})=\bm{K}_{P}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})$ (3)
$diag(k_{v,1},k_{v,2},\cdots,k_{v,n})=\bm{K}_{V}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})$ (4)
where $\bm{\dot{q}}\in\mathbb{R}^{n}$ is the actual velocity of the $n$ actuated joints.
II-A5 The spinal cord module
The spinal cord module combines feedback control coefficients from the brainstem module and the compensation torque from the thalamus module to generate real-time control torque commands.
@@ II-A5 The spinal cord module
In the human CNS, the spinal cord can generate reflexive actions independently via reflex arcs, enabling rapid responses to sensory stimuli without cortical input [28]. Inspired by this mechanism, the spinal cord module produces torque commands at real-time control frequency with a PD-like feedback control mechanism, i.e.,
@@
By integrating spinal feedback control with brainstem-driven adaptive coefficients modulation, the framework effectively addresses non-linearity, Coriolis and centrifugal forces from unknown trajectories.
This reflexive torque is then summed with the compensation torque from the thalamus to generate control torque command $\bm{\tau}$ for actuators, i.e.,
$\bm{\tau}=\bm{\tau}_{\text{fb}}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})+\bm{\tau}_{g}(\bm{q},\bm{F}_{\text{ef}})=(\tau_{1},\cdots,\tau_{n})^{\top}$ (6)
II-B Control levels
Different neural regions in the CNS operate hierarchically during motor control, with each level specializing in particular tasks while communicating and cooperating through neural pathways. This hierarchical organization enables organisms to execute a wide range of motion tasks flexibly and stably in unstructured environments.